(25pt| 3. Find y(t) if x(0) = 0 2 + 2y + u(t), 2.r + y, y(0) = 2 where it is the unit step function.
Answer #1

2+2y + Ult) 110)=0 Y = 22+ Y Y(0)= 2 Так е Laplace + =) L{x} L{x} +2 L{y} L{usces} SL{n} - 210) = L{x} + 2L {y} + (5-1) L{n}s (5-1² -4 £{y} = 0 + 1 S(S-1) a(5-1) + 77 2 (S-1) (S-D²4 sps-18-4 215.1) => [{} 1 Stes=3) (S3_28²35) ai SAI 3_25-25 az s-3 1rake inverse Laplace es 4(5+1) -35 és s-3 - + | e(t-1) + + 3(t-1)/uct-1) Y() = ot [$++e- Y(H) + 3t + + 3(4-1) u(t-1) -t e + e